# LVMark: Robust Watermark for latent video diffusion models

### v2

@@ 6 Conclusion
We proposed LVMark, a first watermarking method for video diffusion models.
Our method captures spatio-temporal information by effectively fusing the 3D wavelet domain and the RGB spatial domain for accurate message extraction from generated videos, even in the case of video modifications. Additionally, our method generates high-quality watermarked videos by selectively modulating latent decoder weights according to the importance of each layer. By balancing the trade-off between video quality and bit accuracy with tailored objective functions, including weighted patch loss, LVMark successfully embedded watermarks into video diffusion models. Experimental results showed that LVMark outperforms all the reasonable baseline methods.
Limitations. Our method demonstrates outstanding performance in the digital watermarking of video diffusion models. However, training the diffusion model to embed the watermark requires substantial memory due to the large size of the video diffusion models and the videos themselves, with memory usage reaching approximately 25GB. Future work could investigate memory-efficient watermarking methods for video diffusion models.

### v3

@@ 5 Conclusion
We proposed LVMark, a robust watermarking method for video diffusion models. Our method captures spatio-temporal information by effectively fusing the 3D wavelet domain and the RGB spatial domain for accurate message extraction from generated videos, even in the case of video and model modifications. Additionally, our method generates high-quality watermarked videos by selectively modulating latent decoder weights according to the importance of each layer. By balancing the trade-off between video quality and bit accuracy with tailored objective functions, including weighted patch loss, LVMark successfully embeds watermarks into video diffusion models. Experimental results show that our method is the first practical video diffusion watermarking approach, achieving a high capacity of 512-bit while preserving original content and ensuring robust watermark retention against various attacks, including model-based ones.
Limitations & Future Work. Our method excels in watermarking video diffusion models. However, training requires significant memory, around 25GB, due to the large size of the diffusion models and videos. Future work could explore memory-efficient watermarking techniques.